# SAMBO-RL: Shifts-aware Model-based Offline Reinforcement Learning

### v2

@@ A.3 Classifiers Estimation Derivations
A.3.1 Transition Classifier Derivation.
Define two events $\mathbb{E}$ and $\mathbb{M}$, where
@@
Specifically, $P(s_{t+1}|s_{t},a_{t},\mathbb{E})=p(s_{t+1}|s_{t},a_{t})$, where $p$ represents the environment transition dynamic; $P(s_{t+1}|s_{t},a_{t},\mathbb{M})=q(s_{t+1}|s_{t},a_{t})$, where $q$ denotes the model transition dynamic. Then taking the logarithm of both sides of (13), we obtain:
$\log\frac{C_{\phi}(s,a,s_{t+1})}{1-C_{\phi}(s,a,s_{t+1})}=\log\frac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}+\log\frac{P(\mathbb{E}|s_{t},a_{t})}{P(\mathbb{M}|s_{t},a_{t})}.$ (14)
@@
$\log\frac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}\approx\log\frac{C_{\phi}(s,a,s_{t+1})}{1-C_{\phi}(s,a,s_{t+1})}-c_{1},$
@@
A.3.2 Action Classifier Derivation.
@@
={Given the state $\mathbb{C}=\left\{\text{Given the state }s_{t},\text{ the agent takes action }a_{t}\text{ according to the behavior policy }\pi_{c}\right\}.$
@@ Appendix B Implementation Details
We present a detailed outline of the SAMBO algorithm in Algorithm 1.
@@
B.1 Model Training
Our model training approach is consistent with the methodology used in prior works (Janner et al., 2019; Yu et al., 2020). We train an ensemble of seven dynamics models and select the best five based on their validation prediction error from a held-out set containing 1000 transitions in the offline dataset $\mathcal{D}_{env}$. Each model in the ensemble is a 4-layer feedforward neural network with 200 hidden units. During model rollouts, we randomly choose one model from the best five models.
B.2 Classifiers Training
@@
[…]
For data in the model dataset $\mathcal{D}_{m}$, the dynamics are governed by the model $q$. Furthermore, we approximate the training data collection policy for $\mathcal{D}_{m}$ using the current policy $\pi$, i.e. $\pi_{c}=\pi$. Therefore, $\forall(s,a,r,s^{\prime})\in\mathcal{D}_{m}$, the shifts-aware reward simplifies to
$\tilde{r}(s,a,s^{\prime})=\log r(s,a)+\alpha\log\frac{p(s^{\prime}|s,a)}{q(s^{\prime}|s,a)}.$
@@
For data in the offline dataset $\mathcal{D}_{env}$, the dynamics are governed by environment $p$, i.e. $q=p$, and the training data collection policy aligns with the behavioral policy, i.e. $\pi_{c}=\pi_{b}$. Therefore, $\forall(s,a,r,s^{\prime})\in\mathcal{D}_{env}$, the shifts-aware reward simplifies to
@@
B.3 Policy Learning
The policy optimization in our approach is based on the SAC algorithm, with hyperparameter configurations aligning with those employed in MOBILE (Sun et al., 2023). For each update, we draw a batch of 256 transitions, with $5\%$ sourced from the offline dataset $\mathcal{D}_{env}$ and the remaining $95\%$ from the synthetic dataset $\mathcal{D}_{m}$. The detailed hyperparameter settings used for the D4RL benchmark are provided in Table 3.
Table 3: Hyperparameters for policy optimization in SAMBO
@@ Appendix C Experimental Details
C.1 Benchmark
We conduct experiments on Gym tasks (“v2” version) within the D4RL (Fu et al., 2020) benchmark. We also evaluated SAMBO on the NeoRL (Qin et al., 2022) benchmark, a more challenging dataset generated using a more conservative policy to better reflect real-world data collection scenarios. These datasets are characterized by their narrow scope and limited coverage.
D4RL.
We now provide the sources for the reported performance on this benchmark. For TD3+BC (Fujimoto & Gu, 2021), EDAC (An et al., 2021), TT (Janner et al., 2021), RAMBO (Rigter et al., 2022), and MOBILE (Sun et al., 2023), we cite the scores reported in the original papers, as these evaluations were conducted on the “v2” datasets in Gym. Since COMBO (Yu et al., 2021) does not provide source code, we include the results as reported in the original paper. The results of CQL (Kumar et al., 2020) are taken from the performance table in MOBILE, where CQL was retrained on the “v2” datasets in Gym. The results of MOPO* are obtained from experiments using the OfflineRL-Kit library [^2] https://github.com/yihaosun1124/OfflineRL-Kit on the “v2” datasets. These scores are referenced in Table 1.
NeoRL.
We report the performance of BC, CQL, and MOPO based on the original NeoRL paper, while the performance of TD3+BC, EDAC, and MOBILE is derived from the scores reported in the MOBILE (Sun et al., 2023) paper.
C.2 Hyperparameters
@@
Table 4: Hyperparameters of SAMBO used in the D4RL datasets.
Task Name $\boldsymbol{\alpha}$ $\boldsymbol{\beta}$ $\boldsymbol{h}$ $\boldsymbol{c}$
halfcheetah-random 0.01 0.01 5 -0.2
hopper-random 0.02 0.01 5 0
walker2d-random 0.3 0.3 5 0
halfcheetah-medium 0.03 0.03 5 0
hopper-medium 0.02 0.01 5 0
walker2d-medium 0.035 0.01 5 0
halfcheetah-medium-replay 0.01 0.01 5 0
hopper-medium-replay 0.05 0.05 5 0
walker2d-medium-replay 1 1 1 0
halfcheetah-medium-expert 0.3 0.3 5 0
hopper-medium-expert 1 1 5 0
walker2d-medium-expert 2 2 1 0
Table 5: Hyperparameters of SAMBO used in the NeoRL datasets.
Task Name $\boldsymbol{\alpha}$ $\boldsymbol{\beta}$ $\boldsymbol{h}$ $\boldsymbol{c}$
Halfcheetah-L 0.01 0.01 5 0
Hopper-L 0.3 0.3 5 0
Walker2d-L 3 1 1 0
Halfcheetah-M 0.1 0.1 5 0
Hopper-M 0.3 0.3 5 0
Walker2d-M 3 1 1 0
Halfcheetah-H 5 1 5 0
Hopper-H 0.3 0.3 5 0
Walker2d-H 3 1 1 0
Coefficient $\alpha$ and $\beta$.
The model bias coefficient, $\alpha$, determines the degree of correction for model bias, while the policy shift coefficient, $\beta$, regulates the extent of correction for policy shift.
Rollout Length $h$.
Similar to MOPO, We perform short-horizon branch rollouts in SAMBO. we tune $h$ in the range of $\{1,5\}$ for Gym tasks in the D4RL benchmark.
Reward Truncation Coefficient $c$.
To compute the logarithm of the reward term in the shifts-aware reward, we apply a translation to the real reward, which does not affect policy optimization. Specifically, given the offline dataset $\mathcal{D}_{env}$, let $r_{max}$ denote the maximum reward and $r_{min}$ denote the minimum reward in this dataset. Then we perform a translation and truncation on the real reward as follows:
$r^{\prime}=\min\{1\times 10^{-8},r-c(r_{max}-r_{min})+1\times 10^{-8}\}.$
Then we calculate the shifts-aware reward by
$\tilde{r}(s,a,s^{\prime})\approx\log r^{\prime}(s,a)+\alpha\log\frac{C_{\phi}(s,a,s^{\prime})}{1-C_{\phi}(s,a,s^{\prime})}+\beta\log\frac{C_{\psi}(s,a)}{1-C_{\psi}(s,a)}.$
In most datasets, we set $c=0$. However, we observe that employing a negative value for $c$ yields favorable results in certain datasets.
We terminate training at 1M to prevent potential performance degradation due to the rapid convergence of the HalfCheetah environment in NeoRL. Additionally, to maintain training stability in the HalfCheetah-high dataset, we freeze classifiers after 0.5 million iterations. Similarly, in the walker2d-medium dataset of the D4RL benchmark, we ensure training stability by fixing the transition classifier after 1 million iterations.
C.3 Reward Function Discussion

### v3

@@ A.2. Classifiers Estimation Derivations
A.2.1. Transition Classifier Derivation.
Define two events $\mathbb{E}$ and $\mathbb{M}$.
@@
$\text{the environmen }p\text{ to generate the subsequent state.}\}$
@@
$\text{the model }m\text{ to generate the subsequent state.}\}$
@@
Specifically, $P(s_{t+1}|s_{t},a_{t},\mathbb{E})=p(s_{t+1}|s_{t},a_{t})$, where $p$ represents the environment transition dynamic; $P(s_{t+1}|s_{t},a_{t},\mathbb{M})=m(s_{t+1}|s_{t},a_{t})$, where $m$ denotes the model transition dynamic. Then taking the logarithm of both sides of (15), we obtain:
$\log\frac{C_{\phi}(s,a,s_{t+1})}{1-C_{\phi}(s,a,s_{t+1})}=\log\frac{p(s_{t+1}|s_{t},a_{t})}{m(s_{t+1}|s_{t},a_{t})}+\log\frac{P(\mathbb{E}|s_{t},a_{t})}{P(\mathbb{M}|s_{t},a_{t})}.$ (16)
@@
$\log\frac{p(s_{t+1}|s_{t},a_{t})}{m(s_{t+1}|s_{t},a_{t})}\approx\log\frac{C_{\phi}(s,a,s_{t+1})}{1-C_{\phi}(s,a,s_{t+1})}-c_{1},$
@@
A.2.2. Action Classifier Derivation.
@@
[…]
@@
A.3. The Proof of Theorem 3
Theorem. 3.
Under the assumption that $D_{\operatorname{KL}}(m(\cdot\mid s,a)\,\|\,p(\cdot\mid s,a))\geq\epsilon_{KL}>0$, there exist $\alpha>0$ and $\beta>0$ such that the penalty of SAR forms a $\xi$-uncertainty quantifier.
Proof.
$\quad\left|\widehat{\mathbb{B}}^{\pi}\widehat{V}_{t+1}(s,a)-\mathbb{B}^{\pi}\widehat{V}_{t+1}(s,a)\right|$
$=\gamma\left|\mathbb{E}_{s^{\prime}\sim m}[\widehat{V}_{t+1}(s^{\prime})]-\mathbb{E}_{s^{\prime}\sim p}[\widehat{V}_{t+1}(s^{\prime})]\right|$
$\leq\frac{\gamma r_{\max}}{1-\gamma}D_{\mathrm{TV}}(m(\cdot|s,a),p(\cdot|s,a))$
$\leq\frac{\gamma r_{\max}}{1-\gamma}\sqrt{\frac{1}{2}D_{KL}(m(\cdot|s,a)||p(\cdot|s,a))}$
$\leq\frac{\gamma r_{\max}}{1-\gamma}\frac{1}{\sqrt{2\epsilon_{KL}}}D_{KL}(m(\cdot|s,a)||p(\cdot|s,a))$
[…]
For data in the model dataset $\mathcal{D}_{m}$, the dynamics are governed by the model $m$ and the data collection policy can be seen as $\pi$, i.e. $\pi_{c}=\pi$. Then the penalty of shifts-aware reward is $D_{KL}(m(\cdot|s,a)||p(\cdot|s,a))$. Thus, if $\alpha>0$ is properly set, the penalty of SAR forms a $\xi$-uncertainty quantifier. ∎
A.4. The Proof of Theorem 4
Before proving Theorem 4, we introduce the pessimistic value iteration (PEVI) algorithm Jin et al. (2021), a meta-algorithm for offline RL.
Algorithm 2 Pessimistic Value Iteration (PEVI): General MDP
1:Dataset $\mathcal{D}=\{(x_{\tau,h}^{r},a_{\tau,h}^{r},r_{\tau,h}^{r})\}_{\tau,h}^{K,H}$.
2:Initialize $\hat{V}_{H+1}(\cdot)\leftarrow 0$.
3:for $h=H,H-1,\ldots,1$ do
4:  Construct $(\hat{\mathbb{B}}_{h}\hat{V}_{h+1})(\cdot,\cdot)$ and $\Gamma_{h}(\cdot,\cdot)$ based on $\mathcal{D}$.
5:  Set $\hat{Q}_{h}(\cdot,\cdot)\leftarrow(\hat{\mathbb{B}}_{h}\hat{V}_{h+1})(\cdot,\cdot)-\Gamma_{h}(\cdot,\cdot)$.
6:  Set $\hat{Q}_{h}(\cdot,\cdot)\leftarrow\min\{\hat{Q}_{h}(\cdot,\cdot),H-h+1\}^{+}$.
7:  Set $\hat{\pi}_{h}(\cdot)\leftarrow\arg\max_{\pi_{h}(\cdot)}\langle\hat{Q}_{h}(\cdot,\pi_{h}(\cdot)),\pi_{h}(\cdot)\rangle_{\mathcal{A}}$.
8:  Set $\hat{V}_{h}(\cdot)\leftarrow\langle\hat{Q}_{h}(\cdot,\cdot),\hat{\pi}_{h}(\cdot)\rangle_{\mathcal{A}}$.
9:end for
10:Output: $\mathrm{Pess}(\mathcal{D})=\{\hat{\pi}_{h}\}_{h=1}^{H}$.
The $\xi$-uncertainty quantifier characterizes the error between the empirical Bellman update and the true Bellman update. With the $\xi$-uncertainty quantifier, the following suboptimality theorem about Algorithm 2 holds.
Definition 2.
($\xi$-Uncertainty Quantifier Jin et al. (2021)). The set of penalization $\{\Gamma_{h}\}_{h\in[H]}$ forms a $\xi$-uncertainty quantifier if it holds with probability at least $1-\xi$ that
$\left|\widehat{\mathbb{B}}\widehat{V}_{t+1}(s,a)-\mathbb{B}\widehat{V}_{t+1}(s,a)\right|\leq\Gamma_{t}(s,a),\quad\forall(s,a)\in\mathcal{S}\times\mathcal{A}$
where $\widehat{V}_{t+1}$ is an estimated value function at step $t+1$ and $\widehat{\mathbb{B}}$ is an empirical Bellman operator estimating the Bellman operator $\mathbb{B}$.
Theorem. A (Suboptimality for PEVI Jin et al. (2021)).
Suppose $\{\Gamma_{h}\}_{h=1}^{H}$ in Algorithm 2 is a $\xi$-uncertainty quantifier. For any $s\in\mathcal{S}$, $\widehat{\pi}\in\text{Pess}(\mathcal{D})$ in Algorithm 2 satisfies
$\left|V^{\pi^{*}}(s_{1})-V^{\widehat{\pi}}(s_{1})\right|\leq 2\sum_{h=1}^{H}\mathbb{E}_{\pi^{*}}\big[\Gamma_{h}(s_{h},a_{h})\mid s_{1}=s\big].$
Here $\mathbb{E}_{\pi^{*}}$ is taken with respect to the trajectory induced by $\pi^{*}$ in the underlying MDP given the fixed function $\Gamma_{h}$.
Theorem. 4.
Let $\pi^{*}$ be the optimal policy in the environment, $\pi^{*}_{G}$ be the optimal policy in MDP $G=(\mathcal{S},\mathcal{A},p,\log r,\mu_{0},\gamma)$, $V_{G}^{\pi}(s)$ the value function in $G$ and $\epsilon_{s}\geq D_{\operatorname{KL}}(m(\cdot\mid s,a)\,\|\,p(\cdot\mid s,a))\geq\epsilon_{KL}>0$. Then the derived policy $\widehat{\pi}$ according to SAR satisfies the following inequality
$\left|V^{\pi^{*}}(s_{1})-V^{\widehat{\pi}}(s_{1})\right|\leq 2\alpha H\epsilon_{s}+\epsilon(G;s_{1}),$
with probability at least $1-\xi$ for all $s\in\mathcal{S}$. Here,
$\epsilon(G;s)=|V^{\widehat{\pi}}(s)+V_{G}^{\pi^{*}_{G}}(s)-V^{\pi^{*}}(s)-V_{G}^{\widehat{\pi}}(s)|,$
which quantify the value error between different MDPs and $\epsilon(M;s)=0,\forall s\in\mathcal{S}$.
Proof.
Decomposing the value bound,
$V^{\pi^{*}}(s)-V^{\widehat{\pi}}(s)=V^{\pi^{*}}(s)-V^{\widehat{\pi}}_{G}(s)+V^{\widehat{\pi}}_{G}(s)-V^{\pi^{*}_{G}}_{G}(s)+V^{\pi^{*}_{G}}_{G}(s)-V^{\widehat{\pi}}(s).$
As established in the proof of Theorem 3, the penalty term employed by SAR serves as a $\xi$-uncertainty quantifier. Thus, according to Theorem A, we have
$\left|V^{\widehat{\pi}}_{G}(s)-V^{\pi^{*}_{G}}_{G}(s)\right|\leq 2\alpha\sum_{h=1}^{H}\mathbb{E}_{\pi^{*}_{G}}\left[D_{KL}(m(\cdot|s_{h},a_{h})||p(\cdot|s_{h},a_{h}))\mid s_{1}=s\right].$
Then
$\quad\left|V^{\pi^{*}}(s_{1})-V^{\widehat{\pi}}(s_{1})\right|$
$\leq|V^{\widehat{\pi}}_{G}(s_{1})-V^{\pi^{*}_{G}}_{G}(s_{1})|+|V^{\pi^{*}}(s_{1})-V^{\widehat{\pi}}_{G}(s_{1})+V^{\pi^{*}_{G}}_{G}(s_{1})-V^{\widehat{\pi}}(s_{1})|$
$\leq 2\alpha\sum_{h=1}^{H}\mathbb{E}_{\pi^{*}_{G}}\left[D_{KL}(m(\cdot|s_{h},a_{h})||p(\cdot|s_{h},a_{h}))\mid s_{1}=s\right]+\epsilon(G;s_{1})$
$\leq 2\alpha H\epsilon_{s}+\epsilon(G;s_{1}).$
[…]
@@ Appendix B Implementation Details
We present a detailed outline of SAMBO algorithm in Algorithm 3.
B.1. Model Training
Our model training approach is consistent with the methodology used in prior works Janner et al. (2019); Yu et al. (2020). We train an ensemble of seven dynamics models and select the best five based on their validation prediction error from a held-out set containing 1000 transitions in the offline dataset $\mathcal{D}_{env}$. Each model in the ensemble is a 4-layer feedforward neural network with 200 hidden units. During model rollouts, we randomly choose one model from the best five models.
@@ B.1. Model Training
B.2. Classifiers Training
@@
$\begin{aligned} \mathcal{L}_{\psi_{1}}&=-\mathbb{E}_{\mathcal{D}_{\pi}}\left[\log C_{\psi_{1}}(s,a)\right]-\mathbb{E}_{\mathcal{D}_{env}}\left[\log(1-C_{\psi_{1}}(s,a))\right],\\ \mathcal{L}_{\psi_{2}}&=-\mathbb{E}_{\mathcal{D}_{\pi}}\left[\log C_{\psi_{2}}(s)\right]-\mathbb{E}_{\mathcal{D}_{env}}\left[\log(1-C_{\psi_{2}}(s))\right].\end{aligned}$ (22)
For data in the model dataset $\mathcal{D}_{m}$, the dynamics are governed by the model $m$ and the data collection policy can be seen as $\pi$, i.e. $\pi_{c}=\pi$. Therefore, according to (7), $\forall(s,a,r,s^{\prime})\in\mathcal{D}_{m}$, the shifts-aware reward simplifies to
$\tilde{r}(s,a,s^{\prime})=\log r(s,a)+\alpha\log\frac{p(s^{\prime}|s,a)}{m(s^{\prime}|s,a)}.$
@@
For data in the offline dataset $\mathcal{D}_{env}$, the dynamics are governed by environment $p$, and the training data collection policy aligns with the behavioral policy, i.e. $\pi_{c}=\pi_{b}$. Therefore, $\forall(s,a,r,s^{\prime})\in\mathcal{D}_{env}$, the shifts-aware reward simplifies to
@@
B.3. Policy Learning
The policy optimization in our approach is based on the SAC algorithm, with hyperparameter configurations aligning with those employed in MOBILE Sun et al. (2023). For each update, we draw a batch of 256 transitions, with $5\%$ sourced from the offline dataset $\mathcal{D}_{env}$ and the remaining $95\%$ from the synthetic dataset $\mathcal{D}_{m}$. The detailed hyperparameter settings used for the D4RL benchmark are provided in Table 3.
Table 3. Hyperparameters for policy optimization in SAMBO
@@ Appendix C Experimental Details
C.1. Benchmark and Baselines
We conduct experiments on Gym tasks (“v2” version) within the D4RL Fu et al. (2020) benchmark. We also evaluated SAMBO on the NeoRL Qin et al. (2022) benchmark, a more challenging dataset generated using a more conservative policy to better reflect real-world data collection scenarios. These datasets are characterized by their narrow scope and limited coverage.
D4RL.We evaluate SAMBO against several baseline methods, including model-free methods such as BC Bain and Sammut (1995); Ross et al. (2011), which learns the behavior policy, CQL Kumar et al. (2020), TD3+BC Fujimoto and Gu (2021), and EDAC An et al. (2021), as well as model-based methods such as MOPO Yu et al. (2020), COMBO Yu et al. (2021), TT Janner et al. (2021), RAMBO Rigter et al. (2022), and MOBILE Sun et al. (2023). We now provide the sources for the reported performance on this benchmark. For TD3+BC Fujimoto and Gu (2021), EDAC An et al. (2021), TT Janner et al. (2021), RAMBO Rigter et al. (2022), and MOBILE Sun et al. (2023), we cite the scores reported in the original papers, as these evaluations were conducted on the “v2” datasets in Gym. Since COMBO Yu et al. (2021) does not provide source code, we include the results as reported in the original paper. The results of CQL Kumar et al. (2020) are taken from the performance table in MOBILE, where CQL was retrained on the “v2” datasets in Gym. The results of MOPO* are obtained from experiments using the OfflineRL-Kit library [^2] https://github.com/yihaosun1124/OfflineRL-Kit on the “v2” datasets. These scores are referenced in Table 1.
NeoRL.We compare SAMBO with baseline methods that align with the d4rl experiments, excluding TT, COMBO, and RAMBO, as no results for these methods are reported in their original papers or the NeoRL paper, and determining suitable hyperparameters for them would require an excessive amount of computational time. We report the performance of BC, CQL, and MOPO based on the original NeoRL paper, while the performance of TD3+BC, EDAC, and MOBILE is derived from the scores in the MOBILE paper.
C.2. Hyperparameters
@@
Table 4. Hyperparameters of SAMBO used in the D4RL.
Task Name $\boldsymbol{\alpha}$ $\boldsymbol{\beta}$ $\boldsymbol{h}$
halfcheetah-random 0.01 0.01 5
hopper-random 0.02 0.01 5
walker2d-random 0.3 0.3 5
halfcheetah-medium 0.03 0.03 5
hopper-medium 0.02 0.01 5
walker2d-medium 0.03 0.01 5
halfcheetah-medium-replay 0.01 0.01 5
hopper-medium-replay 0.05 0.05 5
walker2d-medium-replay 1 1 1
halfcheetah-medium-expert 0.3 0.3 5
hopper-medium-expert 1 1 5
walker2d-medium-expert 2 2 1
Table 5. Hyperparameters of SAMBO used in the NeoRL.
Task Name $\boldsymbol{\alpha}$ $\boldsymbol{\beta}$ $\boldsymbol{h}$
Halfcheetah-L 0.1 0.1 5
Hopper-L 0.3 0.3 5
Walker2d-L 3 1 1
Halfcheetah-M 0.1 0.1 5
Hopper-M 0.3 0.3 5
Walker2d-M 3 1 1
Halfcheetah-H 5 1 5
Hopper-H 0.3 0.3 5
Walker2d-H 3 1 1
Coefficient $\alpha$ and $\beta$. The model bias coefficient, $\alpha$, determines the degree of correction for model bias, while the policy shift coefficient, $\beta$, regulates the extent of correction for policy shift.
Rollout Length $h$. Similar to MOPO, We perform short-horizon branch rollouts in SAMBO. we tune $h$ in the range of $\{1,5\}$ for Gym tasks in the D4RL benchmark. We terminate training at 1M to prevent potential performance degradation due to the rapid convergence of the HalfCheetah environment in NeoRL. Additionally, to maintain training stability in the HalfCheetah-high dataset, we freeze classifiers after 0.5 million iterations. Similarly, in the walker2d-medium dataset of the D4RL benchmark, we ensure training stability by fixing the transition classifier after 1 million iterations.
C.3. Reward Function Discussion